OpenStudy (anonymous):

ALGEBRA 1. MEDAL GIVEN. PLEASE HELP. Part A: Nadia rented a stage platform at $410 for 4 days. If she rents the same stage platform for 8 days, she has to pay a total rent of $770. Write an equation in the standard form to represent the total rent (y) that Nadia has to pay for renting the stage platform for x days. (4 points) Part B: Write the equation obtained in Part A using function notation.(2 points) Part C: Describe the steps to graph the equation obtained above on the coordinate axes. Mention the labels on the axes and the intervals. (4 points)

OpenStudy (anonymous):

for Part A, i got y = (410 / 4)x is that correct???

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

i really only need help with part c.

OpenStudy (anonymous):

b*

zepdrix (zepdrix):

A looks good :) Having trouble with the function notation?

OpenStudy (anonymous):

thank you. :) yes!

zepdrix (zepdrix):

Oh oh maybe we screwed something up... Notice the second piece of information about the 8 days? Hold on hold on hold on >.< gimme a sec

OpenStudy (anonymous):

oh alright. take your time. :)

zepdrix (zepdrix):

Since this is Algebra 1, I'm assuming that we're looking for a linear model. So we want something of the form: \(\Large\rm y=mx+b\)

zepdrix (zepdrix):

We need an equation that will satisfy BOTH pieces of information, not just the first piece.

zepdrix (zepdrix):

Let's take that information and think of it as ordered pairs:\[\Large\rm (4,410)\]\[\Large\rm (8,770)\]

zepdrix (zepdrix):

Understand what I did there? :o

OpenStudy (anonymous):

ummm yes. i believe so.

zepdrix (zepdrix):

The 4 relates to the $410. That's our first set of data. Same with the other stuff. We want to use our slope formula, to figure out the slope of our linear function.

OpenStudy (anonymous):

right

zepdrix (zepdrix):

\[\Large\rm m=\frac{y_2-y_1}{x_2-x_1}\]This look familiar? :o

OpenStudy (anonymous):

yes!

zepdrix (zepdrix):

Plug in the stuff :D Get a number. Give it a try!

OpenStudy (anonymous):

wait, what is y2, y1, x2, x1?? how do i find those??

zepdrix (zepdrix):

Time is independent. Those are what we call our x's. So the number of days are the x's.

zepdrix (zepdrix):

So when I set up my ordered pairs, that's why I put the 4 and 8 in the first slots. Those are what we want to call our x's.

OpenStudy (anonymous):

ohhhhh.

OpenStudy (anonymous):

then the y's would be the cost???

zepdrix (zepdrix):

Yes :) The cost DEPENDS on the amount of days past. That's just something to keep in mind: x - independent y - dependent kinda hard to stuff everything in your brain noggin though :d I know

OpenStudy (anonymous):

so it would be: m = 770 - 410 / 8 - 4 ???

zepdrix (zepdrix):

Mmmm ok the setup looks good! Finish up the calculations to get a single number for your m :)

OpenStudy (anonymous):

m = 90 ???

zepdrix (zepdrix):

\[\Large\rm y=mx+b\]\[\Large\rm y=90x+b\]Ok great! Almost done figuring out our equation.

zepdrix (zepdrix):

So this equation is telling us that it costs 90 dollars PER day to rent the stage. But there might also be some type of "service charge" or something going on. It's a value that doesn't change, it's just always charged to you no matter how many days you rent the stage.

zepdrix (zepdrix):

We'll plug in one of our points to solve for the unknown b.\[\Large\rm (4,410)\qquad\to\qquad ~\quad y=mx+b\]\[\Large\rm (4,410)\qquad\to\qquad 410=90(4)+b\]

OpenStudy (anonymous):

wait, if this equation says it costs 90 dollars per day, my first equation in part A is wrong.

zepdrix (zepdrix):

Yes, we're fixing part A :o

OpenStudy (anonymous):

ohhhhh okay.

zepdrix (zepdrix):

Understand how I plugged that information in? I put the 4 in for the x, the 410 in for the y, then you want to solve for b. \[\Large\rm 410=90(4)+b\]\[\Large\rm 410=360+b\]Understand how to solve for b from this point?

OpenStudy (anonymous):

subtract 360 from 410????

zepdrix (zepdrix):

mmm k good.

zepdrix (zepdrix):

\[\Large\rm 410-360=b\]\[\Large\rm 50=b\]We'll take our b value that we found, and plug it back into our equation. (The one that doesn't have the 4 and 410 plugged in).\[\Large\rm y=90x+b\]\[\Large\rm y=90x+50\]

zepdrix (zepdrix):

Mmmm k so there's part A, yah? :) A lil harder than you were expecting maybe? Hehe

OpenStudy (anonymous):

so part a: y = 90x + 50

zepdrix (zepdrix):

yes.

OpenStudy (anonymous):

okay thank you! haha. could you also help me with part b????

zepdrix (zepdrix):

Part B is nice and easy. Instead of y, we're going to clearly show that the Rent has a dependency on time by writing it as a function of our time variable x.

zepdrix (zepdrix):

So we simply replace y with f(x).

zepdrix (zepdrix):

For part B: \[\Large\rm f(x)=90x+50\]

zepdrix (zepdrix):

And how bout C? You think you can handle the graphing? :)

OpenStudy (anonymous):

thank you so much! and i think...but just in case, could we go over it really quick? :)

zepdrix (zepdrix):

\[\Large\rm y=4x+3\]I'll show you an example :3 and make you do yours on your own hehe

OpenStudy (anonymous):

haha okay.

zepdrix (zepdrix):